The table below displays results from experiments with polygraph instruments. Find the positive predictive value for the test. That​ is, find the probability that the subject​ lied, given that the test yields a positive result.

Did the Subject Actually​ Lie?

No​ (Did Not​ Lie)

Yes​ (Lied)

Positive test results

1515

4040

Negative test results

3030

1212

The probability is?

​(Round to three decimal places as​ needed.)
Answer #1

Solution :

Given that

No​ (Did Not​ Lie) Yes​ (Lied) Total

Positive test results 1515 4040 5555

Negative test results 3030 1212 4242

Total 4545 5252 9797


put A = positive result , B = subject​ lied

=> we know that P(B/A) = P(A and B)/P(A)

= (4040/9797)/(5555/9797)

= 4040/5555

=> P(subject lied/positive result) = 0.727

=> The probability that the subject​ lied, given that the test yields a positive result is 0.727